OpenStudy (anonymous):

what is the prediction equation for the two ordered pairs (5,4635) and (12,5525) ?

OpenStudy (anonymous):

Like a line?

OpenStudy (anonymous):

yeah! i really need some help

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok, you want to find the change in y over the change in x to find the slope of your line. So you'd want to do (5525-4635)/(12-5). This will give you 890/7. Unfortunately that isn't a whole number...but that's your slope. So then you need to plug in a point with that slope to solve y=mx+b for b. Therefore, your line's equation will be y=890/7x + b. To find b solve this equation for b. (I'm sorry I'm being lazy and I don't have paper to write this out and solve it). 4635=(890/7)(5) + b. Once you find b, plug into the above equation with 890/7 as your slope.

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (anonymous):

haha no problem!